Subject: Diffusion Morning Report
From: Ray Seymour <rseymour@stanford.edu>
Date: Fri, 07 Dec 2012 08:21:53 -0800
Fri, 07 Dec 2012 08:21:53 -0800
Good Morning All,
Ted is out today.
*1**. New issues**: PM on Epi 2 LTO N2 alarm
* *2. Ongoing issues:*Tylan Poly rewriting recipes,** Tylannitride  
tested again waiting further instructions after alarms,  working on 
Badger interlocks,*,**O2 in    Silane? TEOS not pumping down, 
**Installing Thermco nitride pump, Epi dome coated-PM today
**  3. Resolved:**

*
					
	Date: 	12/7/12 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*yellow* 	*still a bit thin* 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*14.39/30* 	** 	**
Tylan 8-SiGe 	*down* 	*waiting for users* 	*3.4/30* 	** 	**
Tylan 9-Poly 	*down* 	*rewriting recipes-testing revised recipes* 
*23.7/30* 	** 	**
Tylan 10-Nitride 	*down* 	*tested after sih4 alarm,testing more* 
*28.5/30* 	** 	**
Tylan 11-LTO 	*up* 	** 	*4.3/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*yellow* 	*thin?* 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*up* 	** 	** 	** 	**
Thermco4 	*up* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	*
* 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*down* 	*installing the new pump* 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi Cham. A 	*down* 	*O2 in SiH4? SIMS is back-PM* 	** 	** 	**
AMAT Epi Cham. B 	*down* 	***At the county since 10/30/12* 	** 	** 	**



Good Morning All,
Ted is out today.
  1. New issues: PM on Epi 2 LTO N2 alarm
  2. Ongoing issues:Tylan Poly rewriting recipes,  Tylannitride  tested again waiting further instructions after alarms,  working on Badger interlocks,,O2 in    Silane?  TEOS not pumping down, Installing Thermco nitride pump, Epi dome coated-PM today                                                                                               
  3. Resolved:


                   
  Date: 12/7/12 THICKNESS-Change at    Priority
Tube Name/Number Status        
Tylan 1 up        
Tylan 2 yellow still a bit thin      
Tylan 3 up        
Tylan 4 up        
Tylan 5-BBr3 up        
Tylan 6-POCl3 up        
Tylan 7-TEOS2 up   14.39/30    
Tylan 8-SiGe down waiting for users 3.4/30    
Tylan 9-Poly down rewriting recipes-testing revised recipes 23.7/30    
Tylan 10-Nitride down tested after sih4 alarm,testing more 28.5/30    
Tylan 11-LTO up   4.3/30    
Tylan 12-FGA up        
Thermco 1 yellow thin?      
Thermco 2 up        
Thermco 3 up        
Thermco4 up        
Thermco Poly1 up    
 
Thermco Poly2 up        
Thermco Nitride1 down installing the new pump      
Thermco bpsg up        
AMAT Epi Cham. A down O2 in SiH4? SIMS is back-PM      
AMAT Epi Cham. B down  At the county since 10/30/12